# Radio-frequency cascade readout and coherent exchange control of spin qubits fabricated using a 300 mm wafer process

### v3

@@ Discussion
In conclusion, we have demonstrated exchange control, which forms the basis for two-qubit gates between spin qubits, in an MOS DQD device fabricated using a 300 mm wafer process on natural silicon. The quality of the manufacturing manifests in the low detuning noise and relatively long $T_{2}^{*}$ for a natural silicon MOS device. These results highlight the advantages of industrial manufacturing and encourage follow-up studies in isotopically enriched Si samples, which, during the review of our manuscript, have now been conducted in Ref. [68]. Additionally, building on previous works [21], adapting dedicated gates to primarily control exchange strength over a wider range could (i) enable symmetric exchange pulses and (ii) reduce $J/\Delta E_{z}$ well below 1, which should lead to an overall reduction in sensitivity to charge noise.
Furthermore, we have expanded the portfolio of dispersive readout methods by introducing the radio-frequency electron cascade, which, with its enhanced gain, demonstrates highest performance for dispersive readout in planar MOS devices. Regarding scalability, the radio-frequency driven electron cascade can be extended to larger arrays for distant readout, eliminating the need for swap-based schemes that rely on shuttling information to a sensor [69]. As shown in Fig. 6, this extension builds upon existing schemes for two-dimensional grids using data and ancilla qubits [15, 4]. In this schematic, we assume that the tunnel barriers between each QD can be precisely controlled, enabling optimal readout fidelity through tuning $t_{c}$ [51]. Figure 6(a) illustrates the readout protocol in a simplified one-dimensional array. In order to readout the data qubit QD at the periphery of the unit cell, the data qubit is tunnel-coupled to a neighbouring ancilla qubit QA. The ancilla qubit, in turn, is capacitively coupled to a chain of DQDs configured in the cascade configuration, eventually connected to a reservoir. An rf-drive applied to the ancilla-data DQD initiates the cascade, propagating along the chain to a readout reservoir. This scheme uniquely enables the simultaneous readout of distant data qubits through frequency multiplexing at the readout tank circuit. In contrast to previous work [44], by driving multiple cascade chains at distinct frequencies, multiple distant qubits can be readout simultaneously, as shown in Fig. 6(b). Furthermore, the combination of unit cells into a dense, scalable 2D grid allows readout resource sharing, resulting in the system being resilient to failure points such as faulty QDs or reservoirs, as demonstrated in Fig. 6(c).
@@ Methods
Minimum integration time and readout fidelity calculation. The SNR of the interdot charge transition shown in Fig. 2(b) is determined by fitting the signal to an expression proportional to the quantum capacitance $\propto\partial^{2}E/\partial\varepsilon^{2}$ [42, 40],
@@
$\text{SNR}=\frac{S_{I}^{2}+S_{Q}^{2}}{0.25(\sigma_{I}^{2}+\sigma_{Q}^{2})}$ (6)
where $S_{{\rm rf,I(Q)}}$ is indicated in Fig. 2(b), and $\sigma_{\rm rf,I(Q)}$ is the standard deviation in background signal (away from zero detuning). Here, we assume a white noise spectrum dominated by the first cryogenic amplification stage [70]. In the measurements acquired throughout the text, the applied rf-power is $P_{\text{rf}}=-78~{}$dBm, corresponding to a power $\text{SNR}=1010\pm 40$. The minimum integration time is determined using [40],
@@
where $N_{\text{avg}}=4000$ is the number of averages used in the measurement, and $\Delta f=1.53f_{\text{LPF}}$ is the measurement bandwidth set by the low-pass filter cutoff frequency $f_{\text{LPF}}=100~{}$kHz. Together, these two parameters give the noise equivalent integration time $\tau_{\rm NE}=13~{}{\rm ms}$ for the measurement in Fig. 2(b). Substituting the SNR into eq. (7) we find a minimum integration time $\tau_{\text{min}}=13.0\pm 0.5~{}\upmu$s. In combination with the relaxation time $T_{1}$ extracted in Fig. 2(d), the readout fidelity in eq. (2) can be calculated.
@@
Extended Data Fig. S1: |||| Tuning the radio-frequency driven electron cascade. a-e, Charge stability diagrams as a function of the DQD gate voltages $V_{\text{G}_{1}}-V_{\text{G}_{2}}$. The multi-electron QD bias $V_{G_{\text{S}}}$ is varied in each panel, defined with respect to $V_{\text{G}_{\text{S}}}^{\text{casc}}$, the bias where the conditions for cascade are observed, shown in panel c. This is the bias utilized in the main text of the manuscript. f-j, Simulated charge stability diagrams using the methods described in Supplementary Section S4. k-o Schematic depiction of the electrochemical levels for the corresponding bias conditions, arranged column-wise.
@@
S3 MAGNETOSPECTROSCOPY SIMULATION
$P^{\text{th}}_{i}=\exp(-E_{i}/k_{\text{B}}T_{\text{e}})/Z.$ (20)
S4 CHARGE STABILITY DIAGRAM SIMULATION
@@
where $Z$ represents the partition function, and $\epsilon_{k}$ represents the energy required for quantum dot array to be in the Fock state $\Lambda_{k}$. Substituting this into eq. (26), we obtain;
@@
This term can be further substituted into eq. (25), to obtain;

### v4

@@ Discussion
In conclusion, we have demonstrated exchange control, which forms the basis for two-qubit gates between spin qubits, in an MOS DQD device fabricated using a 300 mm wafer process on natural silicon. The quality of the manufacturing manifests in the low detuning noise and relatively long $T_{2}^{*}$ for a natural silicon MOS device. These results highlight the advantages of industrial manufacturing and encourage follow-up studies in isotopically enriched Si samples, which, during the review of our manuscript, have now been conducted in Ref. [66]. Additionally, similar to previous work [21, 13], adapting dedicated gates to primarily control exchange strength over a wider range could (i) enable symmetric exchange pulses and (ii) reduce $J/\Delta E_{z}$ well below 1, which should lead to an overall reduction in sensitivity to charge noise.
Furthermore, we have expanded the portfolio of dispersive readout methods by introducing the radio-frequency electron cascade, which, with its enhanced gain, demonstrates the highest performance for dispersive readout in planar MOS devices. Regarding scalability, the radio-frequency driven electron cascade can be extended to larger arrays for distant readout, eliminating the need for swap-based schemes that rely on shuttling information to a sensor [67]. As shown in Fig. 6, this extension builds upon existing schemes for two-dimensional grids using data and ancilla qubits [15, 4]. In this schematic, we assume that the tunnel barriers between each QD can be precisely controlled, enabling optimal readout fidelity through tuning $t_{c}$ [51]. Figure 6(a) illustrates the readout protocol in a simplified one-dimensional array. In order to readout the data qubit QD at the periphery of the unit cell, the data qubit is tunnel-coupled to a neighbouring ancilla qubit QA. The ancilla qubit, in turn, is capacitively coupled to a chain of DQDs configured in the cascade configuration, eventually connected to a reservoir. An rf-drive applied to the ancilla-data DQD initiates the cascade, propagating along the chain to a readout reservoir. This scheme uniquely enables the simultaneous readout of distant data qubits through frequency multiplexing at the readout tank circuit. In contrast to previous work [44], by driving multiple cascade chains at distinct frequencies, multiple distant qubits can be readout simultaneously, as shown in Fig. 6(b). Furthermore, the combination of unit cells into a dense, scalable 2D grid allows readout resource sharing, resulting in the system being resilient to failure points such as faulty QDs or reservoirs, as demonstrated in Fig. 6(c).
@@ Methods
Radio-frequency cascade readout performance. The SNR of the interdot charge transition shown in Fig. 2(b) is determined by fitting the signal to an expression proportional to the quantum capacitance $\propto\partial^{2}E/\partial\varepsilon^{2}$ [42, 40],
@@
$\text{SNR}=\frac{S_{I}^{2}+S_{Q}^{2}}{0.5(\sigma_{I}^{2}+\sigma_{Q}^{2})}$ (5)
where $S_{{\rm rf,I(Q)}}$ and $\sigma_{\rm rf,I(Q)}$ are indicated in Fig. 2(b), representing the signal and the standard deviation in background signal (away from zero detuning), respectively. Here, we assume a white noise spectrum dominated by the first cryogenic amplification stage [68]. In the measurements acquired throughout the text, the applied rf-power is $P_{\text{rf}}=-88.5~{}$dBm, corresponding to a power $\text{SNR}=(3.4\pm 0.1)\times 10^{3}$. The minimum integration time is determined using [40],
@@
where $N_{\text{avg}}=4000$ is the number of averages used in the measurement, and $\Delta f=1.53f_{\text{LPF}}$ is the measurement bandwidth set by the low-pass filter cutoff frequency $f_{\text{LPF}}=100~{}$kHz. Together, these two parameters give the noise equivalent integration time $\tau_{\rm NE}=13~{}{\rm ms}$ for the measurement in Fig. 2(b). Substituting the SNR into Eq. (6) we find a minimum integration time $\tau_{\text{min}}=7.6\pm 0.2~{}\upmu$s. In combination with the relaxation time $T_{1}$ extracted in Fig. 2(e), the readout fidelity can be calculated using the following expression [37],
$\mathcal{F}_{r}=\frac{1}{2}\left[1+\text{erf}\left(\sqrt{\frac{\tau_{\text{int}}}{8\tau_{\text{min}}}}\right)\exp\left(-\frac{\tau_{\text{int}}}{2T_{1}}\right)\right]$ (7)
where $\tau_{\rm int}$ is the integration time and $\text{erf}(x)$ is the Gauss error function. The corresponding readout infidelity $1-\mathcal{F}_{r}$ is shown in Fig. 2(f).
We define the detuning fluctuations associated with the rf excitation to be given by [40],
$\delta\varepsilon_{\rm rms}^{\rm rf}=\frac{\alpha_{21}V_{\rm dev}}{\sqrt{2}},$ (8)
where $\alpha_{21}=\alpha_{\rm R,2}-\alpha_{\rm R,1}$ and $V_{\rm dev}$ is the voltage excitation at the reservoir ohmic contact, from input rf signal $V_{\rm in}$ [69],
$V_{\rm dev}=\frac{2C_{C}Q_{L}V_{\rm in}}{C_{C}+C_{P}}$ (9)
where $Q_{L}=100$ is the loaded quality factor of the resonator depicted in Fig. 1(b), and rms $V_{\rm in}=8.4~{}\upmu{\rm V}$ corresponding to input $P_{\rm rf}=-88.5~{}{\rm dBm}$. We then estimate an upper bound to $\alpha_{21}\geq 0.01$ by re-arranging Eq. (1), assuming $\alpha_{\rm R,ME}=0.35\pm 0.15$. The resulting rf-induced fluctuation in detuning is then, $\delta\varepsilon_{\rm rms}^{\rm rf}\leq 2.7\pm 0.6~{}\upmu{\rm eV}$.
@@
Supplementary Fig. S1: |||| Tuning the radio-frequency driven electron cascade. a-e, Charge stability diagrams as a function of the DQD gate voltages $V_{\text{G}_{1}}-V_{\text{G}_{2}}$. The multi-electron QD bias $V_{G_{\text{S}}}$ is varied in each panel, defined with respect to $V_{\text{G}_{\text{S}}}^{\text{casc}}$, the bias where the conditions for cascade are observed, shown in panel c. This is the bias utilized in the main text of the manuscript. f-j, Simulated charge stability diagrams using the methods described in Supplementary Section S3. k-o Schematic depiction of the electrochemical levels for the corresponding bias conditions, arranged column-wise.
@@
S3 CHARGE STABILITY DIAGRAM SIMULATION
@@
where $Z$ represents the partition function, and $\epsilon_{k}$ represents the energy required for quantum dot array to be in the Fock state $\Lambda_{k}$. Substituting this into equation (25), we obtain;
@@
This term can be further substituted into equation (24), to obtain;
@@
S4 MAGNETOSPECTROSCOPY SIMULATION
$P^{\text{th}}_{i}=\exp(-E_{i}/k_{\text{B}}T_{\text{e}})/Z.$ (32)
Reference Platform Resonator $\tau_{\rm int}~{}(\upmu\text{s})$ $\mathcal{F}_{r}~{}(\%)$ $T_{1}~{}(\upmu s)$
Pakkiam et al. 2018 [32] Donor in Si Off-chip SC 300 82.9 620
[…]
Zheng et al. 2019 [34] Si/SiGe On-chip SC 6 98.0 159
[…]
Supplementary Table 1: In-situ dispersive readout performance across different silicon technology platforms. Here resonators are categorised into three-types, on- and off-chip superconducting (SC) resonators and those constructed using surface mounted (SMD) circuit components.
Reference Platform $\delta\varepsilon_{\rm rms}~{}(\upmu\text{eV})$ $\Delta E_{\rm z,rms}~{}({\rm neV})$ $T_{2}^{*}~{}(\upmu\text{s})$ $T_{2}^{\text{echo}}~{}(\upmu\text{s})$ $T_{\rm 2,mag}^{*}~{}(\upmu\text{s})$
Jock et al. 2018 [26] ${}^{28}\text{Si/SiO}_{2}$ 2.0 $\leq$ 0.2a 1.00 8.4 $\geq$20
Connors et al. 2022 [56] Si/SiGe 2.7 $\leq$1.3b 0.10 0.4 $\geq$9
[…]
Supplementary Table 2: Exchange echo reports in silicon quantum devices, where a singlet-triplet qubit is configured to dephase under exchange $J$ and is refocused via the difference in Zeeman energy between the two dots $\Delta E_{z}$ [47] (in a DQD). Noise parameters $\delta\varepsilon_{\rm rms}$ and $\delta\Delta E_{\rm z,rms}$ describe rms fluctuations in detuning $\varepsilon$ and $\Delta E_{z}$, respectively. Note that here, the dephasing time $T_{2}^{*}$ listed does not represent the longest $T_{2}^{*}$ achieved in each reference, but is the reported value at the detuning $\varepsilon_{\text{echo}}$ where the echo was implemented. The echo time is reported as $T_{2}^{\rm echo}$, and this is contrasted against the magnetic limited dephasing time $T_{\rm 2,mag}^{*}=\lim_{\delta\varepsilon_{\rm rms}\to 0}T_{2}^{*}$, a-bdetermined by $\sqrt{2}\hbar/\lim_{\varepsilon\to\infty}T_{2}^{*}$ in aFig. 4(d) [26] and bFig. 2(d) [56].